Logiweb dictionary of prove

Up Help

0 0 prove 1 1 << testMacro(∗) 2 0 Tester1 3 0 Tester2 4 0 Tester3 5 0 Tester4 6 0 Tester5 7 0 Tester6 8 2 ∗ << test∗

The pyk compiler, version 0.grue.20060417+ by Klaus Grue GRD-2006-12-07.UTC:17:24:53.972444 = MJD-54076.TAI:17:25:26.972444 = LGT-4672229126972444e-6